Question:The volume of balloon filled partially with helium is 30 m3 at the earth's surface where the pressure is 76cm(mercury)and the temperature is 27 degree C.If this balloon rise up to a height where the pressure is 7.6cm(mercury)and the temperature is -54 degree C,what will be the volume of the gas there ?
 

Grade:12th pass

1 Answers

Ajay Kumar
6 years ago
Hi Parmanand,
As you know 
PV=nRT or nR=PV/T where nR is constant so we have
P1V/T= P2V2/T2
where V1 = 30m3,   P1 = 76cm,  T1 = 273+27=300°K
V2 = ?  ,   P2 = 7.6cm, T2 = 273+(-54)=219°K
after putting these values in given equation you will have 
V2 = 219m3.
If you still have doubt Pl ask. 
 
Regards,
Ajay Kumar